Find Three consecutive odd integers such that 3 times the largest results in 34 plus twice the smallest
Answer by ikleyn(52873)

.
Let x be the smallest of the three consecutive odd integers under the question.


Then your equation is

3*(x+4) = 34 + 2x.


3x + 12 = 34 + 2x.


=================================>  x = 34 - 12 = 22.


Answer.  The given problem HAS NO solution.

         Such consecutive odd integers do not exist.

And I have a question: for what reason did you post this DEFECTIVE problem to the forum ??????????